OpenStudy (anonymous):

Check my answer? Choose the correct simplification of (4x3y4z5)(5x4y5z3). . 20x12y20z15 9x7y9z8 9x12y20z15 20x7y9z8 - I picked the last one because I'm pretty sure you at the exponents, however I'm not sure if you add the 4x and 5x or multiply them?

OpenStudy (michele_laino):

hint: \[\left( {4{x^3}{y^4}{z^5}} \right)\left( {5{x^4}{y^5}{z^3}} \right) = 4 \times 5\left( {{x^3}{x^4}} \right)\left( {{y^4}{y^5}} \right)\left( {{z^5}{z^3}} \right) = ...?\]

OpenStudy (anonymous):

So i was correct? :D

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

Thank you! :)

OpenStudy (michele_laino):

Thank you! :)